OpenStudy (peridot):

The endpoints of a line segments have the coordinates of (-6, 3, 4) and (4, -1, 2). What is the midpoint of the segment?

OpenStudy (michaelac959):

still need help

OpenStudy (johnweldon1993):

And of course to add onto @Mehek14 , since there are 3 components in the segments we have to add on the 'x' \[\large Midpoint = (\frac{x_1 + x_2}{2} , \frac{y_1 + y_2}{2} , \frac{z_1 + z_2}{2})\]